where does -2/5, and -0.7 go on a number line from -1.0 to 1.0
Answer by stanbon(75887)

where does -2/5, and -0.7 go on a number line from -1.0 to 1.0
----
-2/5 = -4/10
-0.7 = -7/10
---
The order of the numbers from left to right is:
-1.0...-0.7....-2/5....0....1
==============================================
Cheers,
Stan H.
==========